# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

### v3

@@ 3.3 Group-Adaptive Threshold Optimization
and verifies whether $\Delta_{k}\leq\delta_{\mathrm{fair}}$ for each metric $M_{k}$. If any disparity exceeds $\delta_{\mathrm{fair}}$, it is considered unfair. The evaluation for fairness is not included in the loss function because it assesses performance in a binary manner, fair or unfair (e.g., pass or fail), similar to zero-one loss, rather than providing a continuous gradient that can guide optimization.
@@
By iterating this process, the algorithm 1 produces an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that strike a balance between high classification performance and bounded inter-group fairness disparities. This adaptive thresholding method is particularly effective when probability distributions differ significantly across subgroups, as it tailors the decision boundary to each subgroup’s characteristics rather than relying on a single global threshold. Additional details about the algorithm, including learning and implementation details, are included in Appendix F.1.
@@ 4 Experiments
Our experiments aim to assess the performance of FairOPT in comparison to a static threshold and a single universally optimized threshold. For training, we use the RAID and MAGE datasets, and for testing, we utilized SemEval24, MAGE, and Human v. AI 500K dataset.
4.1 Datasets
4.1.1 Train dataset
The RAID dataset (Dugan et al., 2024) encompasses text samples generated by a wide array of language models, including GPT-4, GPT-3, Llama 2 70B, Cohere, MPT-30B, and Mistral 7B. These models produced text across diverse domains such as abstracts, recipes, reddit posts, book summaries, news articles, and poetry. MAGE (Li et al., 2024) includes human-written text covering opinion statements, news articles, question answering, stories, and scientific writing, with machine-generated texts produced by 27 LLMs, including OpenAI’s GPT variants, LLaMA series, FLAN-T5, and EleutherAI models. We randomly extracted 10,000 samples from each of these two datasets.
4.1.2 Test dataset
We also use three datasets for test: SemEval24, MAGE, and Human v. AI 500K. The SemEval24 dataset (Wang et al., 2024) was employed for its diverse representation of commonly used AI text generators balanced with human-written texts across various topics such as news, social media posts, and creative writing, from which we extracted 3,000 samples. We also extracted random 3,000 writings from the MAGE dataset except for the instances that used in the training, and randomly selected 2,000 samples from the Human v. AI 500K dataset, the most widely used dataset for AI text detection on Kaggle.
@@ 4.2 AI-Generated Text Classifiers
To detect AI-generated text, we employed a combination of openly accessible classifiers, RoBERTa-base (Solaiman et al., 2019), RoBERTa-large (Solaiman et al., 2019), and RADAR (Hu et al., 2023). These classifiers were selected based on their open-source availability and the ability to produce unified probabilistic outputs ranging from 0 to 1 with seven decimal points. GPT4o-mini was incorporated for its promising application in AI content detection tasks (Bhattacharjee & Liu, 2024). Unlike binary classification approaches, these models assign a probability score indicating the likelihood that a given text was generated by AI rather than directly making a discrimination. To maintain uniformity in our evaluation, GPT4o-mini was explicitly prompted to return probability scores within the same range as other detectors.
Figure 3: Histograms show how AI-generated probability distributions differ by text length (Short in red, Medium in green, Long in blue). If we use the threshold of 0.5 (black dash line), human-written medium-length text shows higher error rate than other lengths. Each kernel density estimation (KDE) curve reflects the probability scores assigned by RoBERTa-large on the training dataset.
@@ 4.3 Subgroup Attributes
We extracted subgroup attributes from each dataset to undercover the potential disparities present in AI classifiers. These attributes are shown in Table 1 and include both non-sensitive attributes like text length and sensitive attributes inferred from stylistic traits.
@@
Text Length Number of characters in a text Short, medium, long (n = 3)
Stylistic personality Inferred Big Five personality traits. Openness, conscientiousness, extroversion, agreeableness, neuroticism (n = 5)
Table 1: The subgroup attributes studied in our experiments.
Figure 4: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector probabilities. The visualization is based on KDE. The biggest discrepancy is observed with a KS statistic of $0.2078$ ($p<0.01$), while the smallest is $0.1001$ ($p<0.01$). These discrepancies indicate varying levels of divergence between groups based on the characteristics of the given text.
(a) RoBERTa-large
(b) RoBERTa-base
Figure 5: The FairOPT algorithm enables us to obtain a balanced tradeoff between performance and fairness, through a relaxed fairness criterion specified by a given fairness threshold. We improve the Pareto frontier of performance and fairness compared to static and ROCFPR baselines.
Text length measures the number of characters in a text. OpenAI’s evaluations (OpenAI, 2023) revealed performance disparities between short (under 1,000 characters) and long text when using their AI classifier. Building on this and using exploratory data analysis, we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500).
The stylistic personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness (Minej, 2023). We determine the personality trait by selecting the one with the highest probability among the five using a BERT-based personality prediction model (Minej, 2023), which is fine-tuned from BERT-base-uncased (Devlin et al., 2018). Details are noted in Appendix D.
@@ 4.4 Discovery of Subgroup Bias in Text Detection
Our first experimental analysis uncovers many discrepancies in AI-text detection performance with respect to different text lengths and writing styles. For example, as shown in Figure 3, medium-length human-written text is more likely to be flagged as AI-generated than short-length human-written text. Similarly, in Figure 2, long text with a neurotic style is also more likely to be detected as AI-generated.
To obtain a holistic view of these biases, we used the Kolmogorov-Smirnov (KS) test to evaluate distributional discrepancies across various text groups under the assumption of a non-parametric distribution. As shown in Figure 4, there are notable differences between certain subgroup pairs. For instance, the comparison between texts characterized by neuroticism and medium-length versus those with neuroticism and short-length yielded a KS statistic of 0.2078 ($p<0.01$). Extroversion and mid-length and neuroticism and short length, showed a KS statistic of 0.1959 ($p<0.01$). On the other hand, relatively smaller discrepancies were observed between groups such as neuroticism and short length versus openness and short length, with a KS statistic of 0.1155 ($p<0.001$), and between neuroticism and short length versus openness and mid-length, with a KS statistic of 0.1001 ($p<0.01$). These observed discrepancies show that varying subgroup characteristics can lead to diverging outcomes when using a fixed threshold for classification. Addressing these discrepancies through tailored threshold adjustments or other interventions is crucial to ensuring equitable treatment across all subgroups.
Detector Method SemEval Kaggle MAGE
F1 ACC FairLen FairPer FPR F1 ACC FairLen FairPer FPR F1 ACC FairLen FairPer FPR
RoBERTa-large Static 0.4382 0.2847 0.1444 0.1377 0.9892 0.4309 0.2823 0.0607 0.1161 0.9830 0.7147 0.6360 0.1817 0.1760 0.6464
ROCFPR 0.0740 0.2990 0.2197 0.1792 0.4854 0.3469 0.3900 0.1490 0.5782 0.6364 0.5201 0.6475 0.3471 0.2777 0.1031
FairOPT 0.4849 0.3223 0.1347 0.1075 0.9892 0.4542 0.2973 0.0291 0.0755 0.9835 0.7088 0.6080 0.1749 0.1209 0.6473
RADAR Static 0.5592 0.6373 0.3826 0.3399 0.2266 0.7097 0.7373 0.3088 0.2876 0.3360 0.4364 0.3930 0.5979 0.4599 0.6896
ROCFPR 0.0153 0.5293 0.1162 0.0475 0.0019 0.0351 0.6333 0.1199 0.1412 0.0005 0.3215 0.4935 0.4102 0.4811 0.2662
FairOPT 0.6307 0.6620 0.3536 0.4939 0.2228 0.6445 0.6183 0.2269 0.1566 0.3732 0.4541 0.3820 0.5705 0.4140 0.6965
RoBERTa-base Static 0.3345 0.2133 0.0976 0.0978 0.9703 0.3659 0.2457 0.0950 0.1294 0.9553 0.7020 0.6515 0.2023 0.2124 0.5265
ROCFPR 0.0740 0.2253 0.2463 0.1428 0.7633 0.3374 0.2943 0.0876 0.1276 0.8171 0.5814 0.6580 0.4275 0.3666 0.1739
FairOPT 0.3548 0.2253 0.1072 0.0866 0.9696 0.3706 0.2447 0.0876 0.1234 0.9553 0.7122 0.6485 0.1966 0.1517 0.5265
GPT4o-mini Static 0.3737 0.5497 0.1806 0.3137 0.2114 0.2949 0.3083 0.1042 0.2494 0.7390 0.2017 0.4340 0.2921 0.1990 0.2878
ROCFPR 0.0004 0.5247 0.0053 0.0130 0.0057 0.0000 0.6267 0.0005 0.0020 0.0005 0.0000 0.5080 0.0064 0.0039 0.0020
FairOPT 0.4371 0.5553 0.1907 0.6844 0.2139 0.3792 0.3223 0.1279 0.3471 0.7459 0.2802 0.4170 0.2578 0.1697 0.3006
Table 2: Comparison of three thresholding methods (Static, ROCFPR, FairOPT) across four probabilistic AI text classifiers (RoBERTa-large, RoBERTa-base, RADAR, GPT4o-mini) and three data sources (SemEval, Kaggle, MAGE). Metrics include F1, ACC, FairLen (BER discrepancy for Length), FairPer (BER discrepancy for Personality), and FPR. Compared to the Static method, FairOPT results in an average increase of 2.91% in F1 score and decrease of 1.77% in FairPer, while causing a decrease of 0.58% in ACC and an increase of 1.59% in FairLen., demonstrating its effectiveness in enhancing specific aspect of performance and fairness with minimal trade-offs in classifier performance. Although ROCFPR method showed the best result in decreasing the FPR of the model, it overly sacrificed performance.
@@ 4.5 Mitigating Subgroup Bias using FairOPT
Given these findings, we now describe the experimental setup using FairOPT for more robust AI-text detection across subgroups. We extracted the text length and personality features from every item in the dataset, which enabled the partition of 15 subgroups with different characteristics. For the training dataset, the AI text classifiers RoBERTa-base, RoBERTa-large, and RADAR were employed to generate predicted probabilities indicating the likelihood that a given text was AI-generated. These probabilities ranged from 0 to 1, with no direct classification decisions made at this stage. For the test dataset, additional predictions were generated using the GPT4o-mini API.
We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup. We include all hyperparameters and training details in Appendix F.3. Threshold optimization was performed on the training dataset based on the results only from RoBERTa-large rather than averaging the probabilities of three detectors. The relaxed fairness criteria in 3.2 is employed to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness. The RoBERTa-large model was also used to implement the AUROC-based approach (Method based on ROC curve B ) with $FPR<0.1$ constraint for comparison, and the static threshold 2 does not need a model for training in generating a decision threshold.
The optimized thresholds were then applied to the test dataset. Performance was evaluated using F1, ACC, BER, and BER calculated per subgroup. FairLen and FairPer metrics were utilized to quantify the maximum discrepancies in BER across groups defined by text length and personality traits. These metrics are defined as follows:
FairLen $=\max_{a\in L}\text{BER}(a)-\min_{a\in L}\text{BER}(g)$ (9)
FairPer $=\max_{b\in P}\text{BER}(b)-\min_{b\in P}\text{BER}(g)$ (10)
where $L$ includes three groups based on text length, $P$ includes five groups based on big five personality traits, and BER denotes the balanced error rate for each group. A higher value of FairLen or FairPer indicates a greater disparity in classification performance across subgroups.
@@ 4.6 Results with FairOPT for AI-Text Detection
Table 2 presents a comprehensive comparison of three thresholding methodologies—Static, ROCFPR, and FairOPT—applied to four AI text detectors (RoBERTa-base, RoBERTa-large, RADAR, GPT4o-mini) across three distinct datasets (MAGE, Kaggle, SemEval). The evaluation metrics include F1, ACC, FairLen (maximum BER discrepancy for text length), FairPer (maximum BER discrepancy for personality traits), and FPR. Overall across all datasets and tasks tested, using FairOPT yields an average increase of 2.91% in F1 Score and a decrease of 1.77% in FairPer, thereby strictly improving the performance-fairness tradeoff. Overall accuracy was only reduced by a marginal 0.58%, indicating that FairOPT does not affect performance accuracy. These results indicate that FairOPT effectively enhances the overall classification performance and reduces fairness disparities across subgroups, outperforming the Static threshold approach by achieving higher F1 scores and lower BER discrepancies related to stylistic features.
We now analyze the results of FairOPT on each specific dataset. On the SemEval task, overall F1 scores show significant improvement, with RADAR achieving a notable increase of 7% after running our FairOPT algorithm. These findings imply that environments exhibiting patterns similar to SemEval can greatly benefit from this method. Additionally, the RoBERTa-large model exhibits a substantial reduction in discrepancy across all datasets that it was tested on - in the MAGE dataset, the discrepancy related to writing style based on personality decreases by 5%. This implies that models that work similarly to RoBERTa-large can pursue better fairness through our method. For example, RoBERTa-based models also exhibit reduced disparities in text length and style. These results provide strong evidence to support the future application of FairOPT to other models, particularly detectors currently under development, as outlined in OpenAI (2023).
4.7 Tradeoff Analysis between Performance and Fairness
Figure 5 illustrates the accuracy-fairness tradeoffs achieved by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines. The relaxed fairness criterion from Section 3.2 effectively navigates the trade-off space, enhancing the Pareto frontier between classification performance and fairness metrics. Especially, the relaxed fairness criterion leads to early termination of the algorithm before overly sacrificing ACC while still satisfying a reasonable level of fairness. In contrast, the ROCFPR baseline, while achieving lower FPRs as noted in Table 2, significantly compromises TPR thus harming overall ACC, highlighting the superiority of FairOPT in maintaining a harmonious balance between constraints and model performance.
The ROCFPR method demonstrates a notable ability to minimize the FPR across various classifiers and datasets. However, this reduction in FPR comes at a significant cost to the TPR, leading to a sacrifice of ACC as illustrated in Appendix B. This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text.
Conversely, the adaptive thresholding approach employed by FairOPT offers advantages in F1 and group fairness over both static and ROCFPR methods by tailoring decision boundaries to subgroup-specific probability distributions. This strategy ensures that each subgroup, characterized by attributes such as text length and personality traits, is evaluated with an appropriate threshold, thereby balancing the trade-off between fairness and performance. FairOPT achieves this balance by improving F1 scores and reducing FairPer discrepancies without substantially sacrificing ACC, thereby enhancing the classifier’s overall performance and addressing fairness.
@@ 5 Conclusion
This paper proposed FairOPT, based on the observation that there is a statistically significant difference in AI-generated text classifier outputs across subgroups of text (e.g. length and personality) and that this discrepancy can lead to notably higher error rates in specific groups. To address this problem, we proposed FairOPT, a new method for group-specific threshold optimization in AI-generated content classifiers without sacrificing performance. Unlike a universal threshold, which presumes well-calibrated probabilities across varied text distributions, our method effectively captures the variations of the probability distribution across different subgroups, leading to a more robust and fair classification.
While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio). This method will be highly beneficial in post-processing, guiding the model toward enhanced performance or increased fairness, contingent upon the specific requirements and deployment context. Furthermore, while our primary focus has been on improving fairness in the post-training stage, we hypothesize that the proposed method can also be integrated as an in-training fairness enhancement mechanism. Finally, other approaches for threshold optimization, including Bayesian methods and convex optimization, can also be studied in future research.

### v4

@@ 3.3 Group-Adaptive Threshold Optimization
and verifies whether $\Delta_{k}\leq\delta_{\mathrm{fair}}$ for each metric $M_{k}$. If any disparity exceeds $\delta_{\mathrm{fair}}$ (e.g., $0.2$), it is considered unfair. The fairness evaluation is not included in the main loss function because the discrepancy is evaluated in a binary manner: fair or unfair (e.g., pass or fail), similar to zero-one loss.
@@
By iterating this process, the Algorithm 1 converges to an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that reduces the disparity while satisfying the demanded performance criteria. For full implementation details and hyperparameter specifications, refer to Appendix H.1.
Figure 5: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector results in cumulative probability distribution function (CDF; y-axis) based on the training dataset in the range of 0.1 to 0.9 (x-axis). The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$). Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions by stylistic features.
Figure 6: Threshold optimization based on the AUROC-based method under an $FPR\leq 10\%$ constraint (see Appendix C). The ROC curve (blue line) illustrates the trade-off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
@@ 4 Experiments
4.1 Dataset
Our experiments aim to assess the performance of FairOPT in comparison to the static and the AUROC-based threshold optimization methods The dataset is constructed with the RAID [36], MAGE [37], and SemEval24 [38] with an even balance between human and AI text to avoid class imbalance. The generator includes GPT and Text-Davinci variants for their common usage, and the dataset spans a diverse range of writing domains, including social media (e.g., Reddit, ELI5), encyclopedic and academic sources (e.g., Wikipedia, arXiv), instructional and creative content (e.g., WikiHow, poetry, recipes), and reviews and summaries (e.g., Yelp, XSum), ensuring broad coverage of both formal and informal text styles. This experiment employed 19,117 instances for the training and 4,779 samples for the testing. We conducted the featured engineering noted in 4.3 and AI detections noted in 4.2 to extract features and get inferences from classifiers.
@@ 4.2 AI-Generated Text Classifiers
To classify AI-generated text, we employed multiple AI text detectors, DeskLib [39], E5LoRA [40], FastDetectGPT [41], RADAR [13], RoBERTa-base and -large [16], and SimpleAI’s detector [42] for open-source availability. Among closed-source models, GPT4.1-nano and GPT4o-mini [43, 44] have been used for LLM-based text detection [7]. These detectors were selected based on their common application and used in both the train and test datasets. The closed-source text detection is only conducted in the test dataset.
Figure 7: FairOPT decreases the Balanced Error Rate (BER) discrepancy of AI detectors compared to other thresholding methods. FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$). The error bar shows the standard deviation from the mean.
@@ 4.3 Subgroup Attributes
We extracted attributes from each text for sub-grouping to capture heterogeneity in text characteristics. These attributes are shown in Table 1. We categorized text length into three classes based on the number of characters as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500). We also labeled text with stylistic personality based on five traits by selecting the one with the highest probability among the five using a BERT-based personality prediction model [45]. Details are noted in Appendix F. These features were chosen because text length captures structural variation and stylistic personality reflects writing tendencies, both of which critically influence detector stability.
@@
Text length Number of characters in text. Short, Medium, Long (k = 3)
Stylistic traits Inferred Big Five personality traits. Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism (k = 5)
Table 1: Subgroup attributes from the feature engineering.
Figure 8: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness rule that is shown in Figure 4. Our method incurs minimal loss of ACC, achieving $\mu=0.6344\pm 0.1518$ compared to Static ($\mu=0.6376\pm 0.1975$) and ROCFPR ($\mu=0.6249\pm 0.2057$) methods.
@@ 4.4 Discovery of Subgroup Bias in Text Detection
We assess detection performance across text-length subgroups under a static threshold $\theta=0.5$. As illustrated in Figure 3, short human-authored samples incur a substantially higher FPR than mid-length and long texts. This length-dependent disparity underscores the necessity of text length adaptive thresholding to mitigate the discrepancy.
Next, we examine the combined effects of length and stylistic personality (Figure 5). By comparing the empirical cumulative distribution functions (CDFs) of AI-likelihood scores for short, conscientious versus long, neurotic texts, a two-sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix G.1.1 for full KS-test details). The observed disparity in the combination of text length and stylistic personality further highlights the need for adaptive thresholding to guarantee equitable detection performance across heterogeneous text cohorts.
@@ 4.5 Mitigating Subgroup Bias using FairOPT
Given these findings, we now describe the experimental setup for applying FairOPT for AI-text detection with reduced discrepancy across subgroups. We extracted the text length and personality features from every instance in the dataset, partitioning the corpus into 15 subgroups. For the dataset, multiple AI text classifiers were employed to generate predicted probabilities indicating the probability that a given text was AI-generated with $p\in[0,1]$.
We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup and applied them as noted in Figure 2. We include all hyperparameters and training details in Appendix H.3. Threshold optimization was performed on the training dataset based on the results from RoBERTa-large [16] rather than averaging the probabilities of multiple detectors. The relaxed fairness criterion is integrated to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness [46].
We also applied an AUROC-based thresholding method using the RoBERTa-large [16] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC-based optimization procedure [26]. The optimized thresholds were then applied to the test dataset and evaluated based on discrepancy, ACC, and F1 score. Metrics are noted in Appendix B.2.
4.6 Tradeoff Analysis between Performance and Fairness
Figure 4 shows the performance-fairness tradeoff by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines. The relaxed fairness criterion from Section 3.2 to less sacrificing ACC while still satisfying a demanded level of fairness.
The ROCFPR method, shown in Figure 6, demonstrates a notable ability to minimize the FPR in training. However, this reduction in FPR comes at a significant cost to the TPR. This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text, although the overall ACC is not overly sacrificed as shown in Figure 8.
@@ 4.7 Results with FairOPT for AI-Text Detection
Figure 7 demonstrates that group-adaptive thresholding from FairOPT consistently lowers the average discrepancy. Our method decreased the discrepancy to $0.4341$ ($\sigma=0.1927$), showing $12\%$ of reduction compared to the static-threshold baseline ($\mu=0.5512$, $\sigma=0.1758$) and outperforms the AUROC-tuned method ($\mu=0.5595$, $\sigma=0.1582$). Notably, E5LoRA shows the greatest decrease, with its BER discrepancy dropping by nearly half (from $0.63$ to $0.33$), and GPT4o-mini achieved a substantial reduction (from $0.60$ to $0.37$). The Desklib experiences a modest 0.003 reduction. Overall, the FairOPT decreased the discrepancy in all AI text classifiers. The detailed results with standard deviation by the dataset, detector, and thresholding methods are in Appendix LABEL:tab:overall_by_detector_and_dataset.
Figure 8 illustrates that FairOPT incurs only a 0.003 of sacrifice in overall ACC and reduction of the standard deviation by 0.04 while outperforming the AUROC-based method. These findings demonstrate that FairOPT substantially mitigates fairness disparities with negligible ACC tradeoff, offering a practical application. Notably, our method shows great effectiveness to commonly used AI detectors like DeskLib [39], FastDetectGPT [41], RADAR [13], and SimpleAI’s detector [42] by securing ACC over $75\%$ while reducing the discrepancy in all cases.
Moreover, the E5LoRA detector [40] exemplifies the ability of FairOPT to transcend the fairness–performance tradeoff. On the SemEval24 dataset, FairOPT achieves a 0.43 reduction in discrepancy while increasing ACC by 0.23. Similarly, on the MAGE, it delivers a 0.25 decrease in discrepancy coupled with a 0.21 ACC increase. These outcomes highlight the broad applicability of adaptive thresholding for improving both fairness and performance across diverse AI-text detection models, as well as other kinds of probabilistic classifiers.
@@ 5 Conclusion
This paper proposes FairOPT, a post-processing algorithm for probabilistic classifiers that reduces the discrepancy while preserving overall accuracy. Based on the observation that there is a statistically significant difference in AI detector outputs by text characteristics (e.g., length and personality), the need to address the variance has been raised because it can lead to notably higher error rates for specific groups. To address this problem, we proposed a new method for group-specific threshold optimization. Unlike universal thresholding methods, which presume evenly distributed probability regardless of the stylistic features of the text, our method effectively captures the variations of the probability distribution of AI text classifiers across different subgroups, leading to classification with a lower discrepancy.
While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio), including deepfake video and audio detectors. This method will be highly beneficial in post-processing, guiding the model toward decreased discrepancy, contingent upon the specific requirements and deployment context. We can also apply this method to all other probabilistic classifications when subgroup discrepancy matters (e.g., recidivism risk prediction or hiring decisions) by learning tailored decision thresholds that substantially mitigate bias with negligible loss in overall accuracy. Furthermore, while our primary focus has been improving fairness in the post-training stage, we hypothesize that the proposed method can also be integrated as an in-training mechanism. Finally, other approaches for threshold optimization, including Bayesian methods and convex optimization, can also be studied in future research.
Although our experiments focus on text-based detectors, the FairOPT framework is modality-agnostic and readily extends to other domains such as image, video, and audio deepfake detection, where probabilistic outputs require post-processing to ensure equitable performance. More generally, FairOPT applies to any probabilistic classification task with subgroup disparities (e.g., recidivism risk prediction, hiring decisions) by learning tailored decision thresholds that mitigate bias with minimal impact on overall accuracy.
6 Limitations
While FairOPT demonstrably reduces subgroup discrepancy on our English-text benchmarks, it relies on only two features (character-count bins and model-inferred Big Five traits). Other text attributes (e.g., topic, syntax, language) may induce additional biases that our current feature set does not capture. Moreover, the stylistic labels are derived from a downstream BERT-based personality predictor, which itself may introduce cultural bias. We focus on relaxing demographic parity—extensions to other fairness definitions (equal opportunity or equalized odds) may require different optimization objectives or could be fundamentally incompatible.